### v1

# HET/HPF observations of Helium in warm, hot, and ultra-hot Jupiters

Abstract
The near-infrared helium triplet line is a powerful tool for studying atmospheric escape processes of close-in exoplanets, especially irradiated gas giants. Line profile fitting provides direct insight into the mechanisms driving atmospheric mass loss of close-in, Jupiter-sized planets. We present high-resolution transmission spectroscopy results for the helium triplet line of sixteen gas giants (). These observations are part of an extensive helium survey conducted using the Habitable Zone Planet Finder spectrograph on the 10 m Hobby-Eberly Telescope. For the first time, we provide constraints on the helium line for HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. Additionally, we are able to confirm previous robust or tentative detections for HD189733 b, HD209458 b, WASP-52 b, WASP-69 b, and WASP-76 b, and non-detections for HAT-P-3 b, WASP-11 b, WASP-80 b, WASP-127 b, and WASP-177 b. We do not confirm the previous helium narrow-band detection in HAT-P-26 b using high-resolution observations. To identify trends within the population of warm, hot, and ultra-hot Jupiters, we combined our results with available helium studies from the literature. As predicted by theory, we find that warm Jupiters with helium detections orbit K-type stars. However, the helium detections in hot Jupiters are found in low-density planets orbiting F-type stars. There are no helium detections in ultra-hot Jupiters. We compiled a list of 46 irradiated gas giants, but more helium studies are needed to increase the sample and improve our understanding of atmospheric mass loss through helium observations.
show]jaume.miquel@austin.utexas.edu
1 Introduction
Gas giant exoplanets, particularly those on short‑period orbits, are the primary laboratories for atmospheric characterization because their large radii and extended atmospheres produce deep and readily detectable transit signals (S. Seager & D. Deming, 2010). Due to their proximity to the host star, they can reach high equilibrium temperatures (), ranging from warm Jupiters (WJs, 1000 K) through hot Jupiters (HJs, 1000–2000 K) to ultra‑hot Jupiters (UHJs, 2150 K, V. Parmentier et al., 2018; M. Stangret et al., 2022). During a transit, starlight filters through the planetary atmosphere, enabling transmission spectroscopy with high-resolution spectrographs, which probes the composition of the upper atmospheric layers. The combination of large scale heights, high irradiation, and short orbital periods makes warm, hot, and ultra-hot Jupiters particularly favorable for atmospheric studies (I. J. M. Crossfield, 2015; N. Madhusudhan, 2019). These planets have provided the first detections of exoplanetary atomic species, and escaping material (D. Charbonneau et al., 2002; A. Vidal-Madjar et al., 2003).
The atmospheric escape of close-in planets is primarily hydrodynamic, driven by the thermal energy deposited into the exoplanet’s atmosphere (R. T. Pierrehumbert, 2010; T. T. Koskinen et al., 2014). Photo-evaporation is the typical hydrodynamic escape process and is driven by stellar irradiation in the most energetic wavelengths, i.e., the X-ray and extreme ultraviolet (XUV111We define XUV as wavelengths up to 504 Å since this range is relevant to helium studies due to the ionization wavelength cutoff of the He i. We note that other works may define the XUV range different, e.g. up to the H ionization cutoff at 912Å.) range (J. E. Owen & A. P. Jackson, 2012). This energy input can inflate the upper atmosphere carrying light elements, such as hydrogen and helium, to higher altitudes. Additionally, the outer layers of the atmosphere of extremely close-in planets can extend beyond the Roche lobe, resulting in a high atmospheric mass loss rate (J. E. Owen, 2019). Roche lobe overflow escape is independent of stellar flux and can act alongside other evaporation mechanisms (A. Lecavelier des Etangs et al., 2004). In extreme cases, these extended, escaping atmospheres can result in the formation of comet‑like tails that interact with the stellar wind, detectable via atmospheric escape tracers, such as the He i triplet line (e.g., Z. Zhang et al., 2023; M. Gully-Santiago et al., 2024). Over time, these mass‑loss processes can sculpt planetary demographics, contributing to observed features such as the sub‑Jovian desert, and providing empirical constraints for models of photo-evaporation (J. E. Owen & A. P. Jackson, 2012; J. E. Owen & D. Lai, 2018).
The He i triplet at 10833 Å has become a key diagnostic for probing the upper atmospheres and escape processes of close‑in exoplanets, including irradiated gas giants (S. Seager & D. D. Sasselov, 2000; A. Oklopčić & C. M. Hirata, 2018; M. Lampón et al., 2021). The absorption arises from the metastable 23S state of neutral helium, which is populated when stellar XUV radiation ionizes helium atoms, and subsequent recombination fills the triplet state. However, the helium metastable state has a lifetime of 2.2 h (G. W. Drake, 1971). Thus, a constant and intense irradiation is needed to sustain detectability of the He i triplet line (J. Sanz-Forcada & A. K. Dupree, 2008). Since its first detection in WJs (J. J. Spake et al., 2018, WASP-107 b; L. Nortmann et al., 2018, WASP-69 b), helium absorption has been reported in various types of exoplanets, providing a direct tracer of atmospheric escape accessible from ground-based telescopes with high‑resolution infrared spectrographs (J. Orell-Miquel et al., 2024). Moreover, ground-based high-resolution spectroscopy observations allow to retrieve several physical parameters from line profile fitting to study the hydrodynamical escape (e.g. M. Lampón et al., 2023; J. Orell-Miquel et al., 2023). Therefore, helium spectroscopy provides a unique window into both the instantaneous escape dynamics and the long‑term sculpting of exoplanet populations.
In this context, this work presents the results of an observational study aimed at characterizing atmospheric escape in gas giants. We have been carrying out a survey targeting the He i triplet line for planets with 0.5 RJup and ranging 800–2100 K around F-, G-, and K-type host stars. Figure 1 presents the planets observed by our survey. The individual analyses and discussions for HAT-P-32 b and HAT-P-67 b are presented in Z. Zhang et al. (2023), and M. Gully-Santiago et al. (2024), respectively. Here, we only included those results, and not new analysis on these two systems in our helium detections discussion.
2 Observations and data analysis
2.1 HPF spectroscopic observations
We obtained spectroscopic observations of gas giant host stars with the Habitable Zone Planet Finder (HPF; S. Mahadevan et al. 2012, 2014; A. J. Metcalf et al. 2019) mounted on the 10 m Hobby-Eberly Telescope (HET; L. W. Ramsey et al. 1998; M. Shetrone et al. 2007). HPF spectrograph covers the near-infrared wavelength range from 0.81–1.38 m with a resolving power of 55 000. HPF data were reduced using the Goldilocks222https://github.com/grzeimann/Goldilocks_Documentation pipeline (J. P. Ninan et al., 2018; K. F. Kaplan et al., 2019), obtaining the science, blank sky, and calibration. To avoid saturation or scattered light from the laser frequency comb, we did not use a simultaneous calibration source in the calibration fiber.
The HET telescope structure can rotate in azimuth but it is fixed at an elevation of 55 degrees. Target tracking is performed by a prime focus instrument package that moves across the focal plane to follow the target. Due to the design of the HET, we are limited to observations during “tracks” of typically 1 hour before (east track) and after (west track) target culmination, making it difficult observe a full transit, or to obtain in-transit and out-of-transit spectra on the same night. Out-of-transit phases are usually covered by observations during the nights surrounding each transit event. HPF observations were organized into visit campaigns around an exoplanet transit. We also obtained random-phase spectra, referred to as P4 observations, to monitor long-term stellar variability. P4 observations are typically used to fill the HET queue or when the conditions are not ideal for higher-priority targets.
We provide a summary of the observations performed for each target analyzed in this work. We report the signal-to-noise ratio (S/N) per extracted pixel of the échelle order where the He triplet line is located.
HAT-P-3 b: We scheduled one visit, centered on the transit on 1 February 2021, during which we collected 16 spectra: 8 in-transit and 8 out-of-transit; complemented with 4 P4 spectra. We set the exposure time to 630 s, achieving a median S/N of 94.
HAT-P-12 b: We collected 21 spectra on visit 1 (31 March 2020): 5 in-transit and 16 out-of-transit; and 12 spectra on visit 2 (27 May 2025): 8 in-transit and 4 out-of-transit. We also collected 2 P4 spectra. We set the exposure time to 960 s for visit 1 and 620 s for visit 2, achieving a median S/N of 52 in both visits.
HAT-P-17 b: We collected 23 spectra on visit 1 (8 June 2020): 4 in-transit and 19 out-of-transit; and 32 spectra on visit 2 (30 July 2020): 7 in-transit and 25 out-of-transit. We also collected 8 P4 spectra. We set the exposure time to 310 s for both visits, achieving a median S/N of 78 and 117, respectively.
HAT-P-26 b: We scheduled one visit, centered on the transit on 1 February 2021, during which we collected 15 spectra: 5 in-transit and 10 out-of-transit; complemented with 2 P4 spectra. We set the exposure time to 660 s, achieving a median S/N of 89.
HD118203 b: We collected 26 spectra on visit 1 (7 January 2021): 14 in-transit and 12 out-of-transit; and 29 spectra on visit 2 (19 February 2021): 18 in-transit and 11 out-of-transit. We set the exposure time to 310 s for both visits, achieving a median S/N of 325.
HD189733 b: We obtained 9, 9, and 2 in-transit spectra for the transits on 4 May 2019, 13 June 2019, and 1 August 2020, respectively. We only collected 11 spectra out-of-transit on the night of 17 April 2019, complemented with 10 P4 spectra across 2020. We set the exposure time to 310 s, achieving a typical median S/N of 400.
HD209458 b: We scheduled one visit, centered on the transit on 2 October 2019, during which we collected 12 spectra: 5 in-transit and 7 out-of-transit. We set the exposure time to 520 s, achieving a median S/N of 310.
TrES-1 b: We scheduled an observational campaign centered on the transits on 7 and 13 July 2020. In total, we collected 33 spectra: 9 in-transit and 24 out-of-transit; complemented with 9 P4 spectra. We set the exposure time to 820 s, achieving a median S/N of 97.
WASP-11 b/HAT-P-10 b: We collected 18 spectra on visit 1 (7 December 2019): 5 in-transit and 13 out-of-transit; and 12 spectra on visit 2 (18 December 2019): 7 in-transit and 5 out-of-transit. We also collected 13 P4 spectra. We set the exposure time to 820 s for both visits, achieving a median S/N of 105.
WASP-52 b: We collected 9 spectra on visit 1 (13 August 2019): 3 in-transit and 6 out-of-transit; and 12 spectra on visit 2 (10 October 2019): 5 in-transit and 7 out-of-transit. We also collected 13 P4 spectra. We set the exposure time to 820 s for visit 1 and 760 s for visit 2, achieving a median S/N of 85.
WASP-69 b: We collected one useful in-transit spectrum on 20 September 2019, and 23 spectra on visit 2 (21 October 2019): 7 in-transit and 16 out-of-transit. We set the exposure time between 250–500 s, achieving a median S/N of 215.
WASP-76 b: We collected 22 spectra on visit 1 (2 October 2020): 10 in-transit and 12 out-of-transit; and 18 spectra on visit 2 (4 October 2020): 10 in-transit and 8 out-of-transit. We set the exposure time to 320 s for both visits, achieving a median S/N of 135.
WASP-80 b: We scheduled one visit, centered on the transit on 13 July 2019, during which we collected 11 spectra: 8 in-transit and 3 out-of-transit; complemented with 16 P4 spectra. We set the exposure time to 310 s, achieving a median S/N of 125.
WASP-127 b: We scheduled one visit, centered on the transit on 13 March 2019, during which we collected 35 spectra: 10 in-transit and 25 out-of-transit; complemented with 3 P4 spectra. We set the exposure time to 415 s, achieving a median S/N of 123.
WASP-156 b: We collected 14 spectra on visit 1 (24 October 2020): 6 in-transit and 8 out-of-transit; and 18 spectra on visit 2 (20 November 2020): 5 in-transit and 13 out-of-transit. We set the exposure time to 590 s for visit 2 and 550 s for visit 2, achieving a median S/N of 70 and 100, respectively.
WASP-177 b: We scheduled an observational campaign centered on the transits on 20 and 23 August 2020. In total, we collected 35 spectra: 8 in-transit and 27 out-of-transit. We set the exposure time to 830 s, achieving a median S/N of 85.
2.2 Telluric correction
The main objective of these observations was the analysis of the planetary He i triplet line, which is surrounded by H2O absorption and OH emission lines from the Earth’s atmosphere. The relative positions of the He i triplet and telluric spectral lines change with the epoch due to the Doppler shift induced by the barycentric Earth radial velocity. Thus, we scheduled the campaigns at epochs that minimize the overlap of the He i planetary trace and the telluric lines (J. Orell-Miquel et al., 2022). Moreover, we also correct the science spectra from telluric lines to minimize their impact on the planetary signals. We use the python package muler (M. Gully-Santiago et al., 2022) to correct the target spectra from the telluric emission lines. The muler code uses the simultaneous sky spectra and accounts for the different efficiencies between target and sky fibers (Z. Zhang et al., 2023; M. Gully-Santiago et al., 2024). We use the telluric model from Goldilocks pipeline, which accounts for the instrumental profile and spectral resolution, to empirically fit and remove the telluric absorption lines (F. Yan et al., 2015; N. Casasayas-Barris et al., 2017).
2.3 Transmission spectrum analysis
We analyzed the HPF spectroscopic time series using the single line transmission spectroscopy technique, which compares the in-transit spectra with the out-of-transit spectra. We adapted the methodology presented in J. Orell-Miquel et al. (2022, 2023, 2024) to the particularities of the HET following Z. Zhang et al. (2023) and M. Gully-Santiago et al. (2024). Table 3 shows the planetary and stellar parameters needed to compute the transmission spectrum for our targets.
Once corrected from the telluric contributions, we normalized each spectrum by its own continuum. We used muler to fit a first order Chebyshev’s polynomial to featureless spectral regions. Then, we shifted the normalized spectra into the stellar rest frame. We computed a high S/N stellar spectrum from the out-of-transit data (MasterOut). To minimize the impact of stellar variability, we analyzed each in-transit observation individually. For each transit visit, we computed a different MasterOut spectrum using only the out-of-transit spectra that are closer to the in-transit track. We divided all spectra (in-, out-of-transit, and P4) by the MasterOut to remove the stellar contribution and to obtain the residual map in the stellar rest frame. At this step, we computed the stellar light curve of the He i triplet integrating the counts within a 1.5 Å band pass centered at the doublet position. The light curves are shown in the Figures 13 and 14, and provide information of the stellar variability and the presence of large planetary tails (Z. Zhang et al., 2023; M. Gully-Santiago et al., 2024).
We moved the residual map into the planet rest frame accounting for the planet movement around the star. We neglected the planets’ eccentricity because it is usually poorly constrained and the impact is within the uncertainties derived for the signal shifts, except for HD118203 b, HAT-P-17 b, and HAT-P-26 b. For these three planets, we used the radvel (B. J. Fulton et al., 2018) package to compute the radial velocity contribution of the eccentricity during the transit. We computed the transmission spectrum (TS) by averaging the in-transit spectra weighted by the inverse of the squared propagated errors. If we obtained more than one in-transit visit, we computed the final TS from the combined residual maps. When the TS shows an absorption feature, we fitted the signal with a Gaussian profile using the Markov’s chain Monte Carlo algorithm via its python implementation emcee (D. Foreman-Mackey et al., 2013). If there were no clear signals in the TS, we derived conservative limits for the planetary absorption. We computed a 3 upper limit as three times the root-mean-squared value of a flat spectral region of the TS.
3 Transmission spectroscopy results
| Planet | Depth | EW | FWHM | Singlet depth | ||
| (%) | [mÅ] | [km s-1] | [Å] | [Å] | (%) | |
| HD189733 b | 0.410.08 | 1.880.35 | 1.71.2 | 0.180.03 | 0.430.07 | – |
| HD209458 b | 0.670.08 | 6.26 | 6.51.5 | 0.37 | 0.88 | – |
| WASP-52 b | 1.850.20 | 37.9 | 3.12.5 | 0.82 | 1.93 | – |
| WASP-69 b | 3.500.10 | 43.41.1 | 5.60.4 | 0.4600.017 | 1.080.04 | 0.680.12 |
| WASP-76 b | 0.480.09 | 6.01.0 | 2.7 | 0.48 | 1.15 | – |
| Planet | Depth () | EW [mÅ] |
| HAT-P-3 b | 1.2 | 12.7 |
| HAT-P-12 b | 2 | 21.2 |
| HAT-P-17 b | 0.7 | 7.4 |
| HAT-P-26 b | 1.2 | 12.7 |
| HD118203 b | 0.2 | 2.1 |
| TrES-1 b | 1.1 | 11.6 |
| WASP-11 b | 1.2 | 12.7 |
| WASP-80 b | 0.8 | 8.5 |
| WASP-127 b | 0.5 | 5.3 |
| WASP-156 b | 0.9 | 9.5 |
| WASP-177 b | 1.1 | 11.6 |
In this section, we present the He i triplet results obtained by applying the methodology described in the previous section. The final residual maps and TS for each planet are shown in Figures 2–5. Table 1 reports the results of our Gaussian profile fitting for the helium signals. In Table 2, we report the derived upper limits for planetary He i absorptions from our HPF TS.
Figures 9–12 show the performance of the telluric correction and the comparison between the MasterOut and the individual in-transit spectra. For multiple visits, we also present the individual TS (see Figs 10, 11, and 12). The He i triplet stellar light curves are shown in Figures 13 and 14.
The main objective of this work is the study of the He i triplet line, but other strong spectral features are accessible to HPF. In Appendix B, we provide a note on the study of the Ca ii infrared triplet lines with HET/HPF.
3.1 HAT-P-3 b
The He i triplet stellar line partially overlaps with the OH lines in the only visit to the target, but the telluric emission correction is well performed (Fig. 9). The residual map and TS do not indicate the presence of a planetary absorption in our data. We derive a 3 upper limit to the He i excess absorption of 1.2 %.
3.2 HAT-P-12 b
The planetary trace is slightly overlapped by OH emission lines in the two visits (Fig. 10). While the TS from the first visit shows a broad 2 % excess absorption feature, the TS from the second visit is dominated by noise without a clear absorption signal. When combining both observations, final TS (Fig. 2) looks similar to that of the first visit. The shape of the absorption feature does not appear to be of planetary origin, but rather from residual OH emissions. Although, the S/N of the final TS would be sensitive to a He i triplet excess absorption of 1.3 %, we report the 2 % absorption from the signal at the He i triplet position as upper limit.
3.3 HAT-P-17 b
The stellar line does not overlap with the telluric lines, but the planetary trace is in between the OH emission line and the H2O absorption line due to the RV shift caused by the planet’s eccentricity (Fig. 10). We obtained two visits which sampled the terminators of HAT-P-17 b (visit 1: ingress, visit 2: egress). Although we did not get full in-transit spectra, these observations would still be able to detect evaporation tails. Despite the telluric contamination, the TS are mainly flat without significant features (Fig. 10). The combined TS (Fig. 2) does not present absorption features from the planet atmosphere. We derive a 3 upper limit to the He i excess absorption of 0.7 %.
3.4 HAT-P-26 b
The He i stellar line is free from tellurics, but the planetary trace falls over the blue doublet of the OH emission lines when accounting for the RV shift caused by the planet’s eccentricity (Fig. 9). Despite the telluric overlapping, the TS is mainly flat (Fig. 2) and does not suggest He i excess absorption during the transit. The 3 upper limit to the He i excess absorption in our observations is of 1.2 %.
3.5 HD118203 b
The He i stellar line is not contaminated by telluric lines (Fig. 10). However, the planetary signal falls close to the OH peak at 10835 Å when accounting for the RV shift caused by the planet’s eccentricity. Visit 1 does not show evidence of planetary absorption, while visit 2 presents a broad 0.3 % absorption feature. However, this signal is similar to other features in both TS. The residual map (see Fig. 3) does not show a clear planetary trace and the combined TS is mainly flat. Although the stellar LC is consistently negative at in-transit phases (Fig. 13), we do not consider that the signal is of a planetary origin. If we combine the two visits, we can derive a 3 upper limit to the He i excess absorption of 0.2 %.
3.6 HD189733 b
We obtained three visits that sample almost all the planetary transit (visits 1, 2, and 3 in Fig. 13). The planetary trace does not overlap with the telluric lines (Fig. 9), but the out-of-transit baseline (visit 4 in Fig. 13) is more than one month apart from the in-transit visits. The He i stellar line shows some variability that prevents us from computing the planetary TS. Fortunately, two spectra from P4 observations were taken just a few days after visit 1. We computed the residual map and TS shown in Fig. 3 using those two P4 spectra as out-of-transit and only visit 1 as in-transit. We masked some strong telluric residuals. The TS show a significant absorption feature at the He i triplet position (0.410.08 %). However, we emphasize that there are other features in the TS of the same amplitude.
3.7 HD209458 b
We observed one in-transit visit, where the stellar line is free from tellurics (Fig. 9). Since we detect some stellar variability between visits (Fig. 13), we used the closer out-of-transit visit which has only one good S/N spectrum. Although there are some telluric residuals, the residual map and the TS present a tentative absorption feature at the planetary trace position. If we fit the signal, we derive a blue-shifted excess absorption of 0.670.08 %. However, we note that the amplitude of this signal is similar to that of the stellar variability (Fig. 13).
3.8 TrES-1 b
The stellar line overlaps with the telluric blue OH doublet, but there are no significant telluric residuals in the individual TS (Fig. 10). The residual map and combined TS do not present evidences of planetary absorption (Fig. 3) and we can only derive a 3 upper limit to the He i absorption of 1.1 %.
3.9 WASP-11 b / HAT-P-10 b
The stellar line overlaps with the OH line and some telluric residuals might be present in the individual TS of both visits (Fig. 11). The first visit shows a feature which we fit as excess absorption of 1.00.2 % but with no clear origin. The second visit shows features which seem to have a clear origin in the telluric residuals. The combined residual map (Fig. 4) presents a tentative absorption aligned with the planetary trace, and the TS shows an absorption feature similar to those detected in the individual TS. However, the combined transit is more noisy, influenced by the second transit data. We note that the stellar LC presents variability of similar amplitude to the observed feature (Fig. 14). Therefore, we consider the absorption from visit 1 as the upper limit of the He triplet absorption. The He i detection would need to be confirmed by completely uncontaminated transit observations.
3.10 WASP-52 b
The first visit is slightly contaminated by OH emission, but the TS presents an absorption at the He i triplet position (Fig. 11). We fitted the feature including an offset since the TS around the He i line is consistently below 0 (offset = 0.550.13 %). We derive a blue-shifted (=73 km s-1) excess absorption of 1.70.4 %. The TS of the second visit shows a clear absorption feature as well. In the second case, we fitted an excess absorption of 1.90.3 % (=13 km s-1). Figure 4 shows the residual map and the TS when combining both visits. Our observations provide a clear He i triplet detection with an excess absorption of 1.850.20 %.
3.11 WASP-69 b
The stellar line from both visits slightly overlap with the OH emission line, but there are no obvious telluric residuals. Although the first visit only has one in-transit spectrum with a useful S/N, the He line is clearly deeper than the MasterOut spectrum (Fig. 11). From visit 1, we obtained a noisy TS showing a feature absorption of 52 %. The TS from visit 2 clearly shows an excess absorption signal from the planetary atmosphere. For this planet, we used two Gaussian profiles to fit simultaneously the doublet and singlet lines of the He i triplet. We obtained an excess absorption of 3.500.10 for the doublet and of 0.680.12 for the singlet. The planetary signal presents a significant blue-shift (=5.60.4 km s-1). We report as our final results those from visit 2 alone (Fig. 4).
3.12 WASP-76 b
The He i stellar line does not overlap with telluric lines, although the planetary trace does due to the large velocity shift during the transit (Fig 11). However, we do not detect residuals from the telluric correction. The TS of the first visit shows an absorption feature (0.550.15 %) coincident with the planetary position. The TS of the second night is a bit more noisy with an absorption feature of 0.50.1 %. Moreover, the He i line curve is mainly negative at in-transit phases (Fig. 14). When combining the spectra from both visits (Fig. 4), the final TS shows a clear absorption feature with a significant absorption of 0.480.09 % (=32 km s-1).
3.13 WASP-80 b
3.14 WASP-127 b
3.15 WASP-156 b
The stellar line partially overlaps with the telluric lines, mainly in the second visit (Fig. 12), but the individual TS are mainly flat without residuals from the telluric correction. The residual map and combined TS do not present evidence of planetary absorption (Fig. 5). We derive a 3 upper limit to the He i absorption of 0.9 %.
3.16 WASP-177 b
We obtained two visits with in- and out-of-transit observations in the same night, as well as out-of-transit observations in the surrounding nights. In both visits, the planetary trace does not overlap with telluric lines (Fig. 12). When analyzing visit 1 using the out-of-transit spectra from the surrounding nights, the TS shows a clear absorption-like feature. However, when using the out-of-transit spectra from the same night, this feature seems noisier. We decide to use this latter option (Fig. 12) because it minimizes any stellar variability impact. For visit 2, we found no significant differences when using the out-of-transit spectra from the same night or from one night before and after. The TS from visit 2 is mainly flat, which increases the doubts about the planetary origin of the feature from visit 1. The combined TS (Fig. 5) shows a narrow absorption due to the feature reported in the TS from visit 1. The combined residual map and TS do not provide strong evidence for a planetary absorption. We report a 3 upper limit of 1.1 %.
4 Discussion
4.1 Comparison of HPF results with literature measurements
We analyzed the He triplet line of five planets for the first time: HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. Although we do not firmly detect excess helium absorption, we provide the first constraints on excess absorption for these planets (see Table 2).
We also report the first He i triplet line spectrographic results of HAT-P-26 b, which has previously only been studied with helium narrow-band photometry by S. Vissapragada et al. (2022). We retrieve a flat TS that suggests no He i absorption, and so we set our upper limit on excess absorption at 1.2 %. However, S. Vissapragada et al. (2022) found a significant He i absorption signal with a mid-transit excess depth of 0.31 % using a 6.35 Å FWHM band-pass. We computed the planet’s helium light curve binning the TS to match the FWHM. We measure a 0.1 % excess depth, which is not consistent with the narrow-band results. The scatter present in S. Vissapragada et al. (2022)’s light curves may help to explain the discrepancy between the two observations. This discrepancy could be due to temporal planetary variability, though more observations are needed to explore this hypothesis.
Additionally, we observed ten other planets, all of which have been the subject of previous spectrographic observations. For HAT-P-3 b, we set an upper limit on He i excess absorption of 1.2 %, which improves the constraint of 1.9 % set by G. Guilluy et al. (2023) using GIANO-B. Similarly, we found an upper limit of 1 % for WASP-11 b. R. Allart et al. (2023) previously set the upper limit at 1.56 % with SPIRou for this planet. While we improve previous constraints, additional observations are needed to confirm the presence of metastable helium in either planet’s atmosphere.
M. Salz et al. (2016) first identified WASP-80 b as a strong candidate for atmospheric escape, particularly due to the planet’s high-energy irradiation from its host star (G. W. King et al., 2018). However, past observations by L. Fossati et al. (2022, GIANO-B), R. Allart et al. (2023, SPIRou), and A. Masson et al. (2024, SPIRou) resulted in tentative detections with upper limits of 0.7 %, 1.24 %, and 2.4 %, respectively. Our upper limit agrees with past observations and suggests that WASP-80 b does not have currently detectable helium outflow despite predictions. L. Fossati et al. (2022) considered that a sub-solar He/H ratio in combination with a high stellar wind might contribute to these non-detections. These non-detections illustrate our limited understanding of the factors that drive the development of the He i triplet.
WASP-127 b has a large radius (1.31 RJup), which indicates that the planet has an extended atmosphere and is a good candidate for detecting outflowing helium (K. W. F. Lam et al., 2017). However, spectroscopic observations have only reported upper limits between 0.48-0.9 % (L. A. dos Santos et al., 2020, Gemini/Phoenix; R. Allart et al., 2023, SPIRou; A. Masson et al., 2024, SPIRou). From our HPF TS, we also put an excess absorption upper limit of 0.5 %, which is consistent with the previous results. As L. A. dos Santos et al., 2020 suggested, the metastable He state may not form due to the host star’s old age (11 Gyr, K. W. F. Lam et al., 2017) and low amount of ionizing radiation.
Similarly, J. Kirk et al. (2020) identified WASP-177 b as a good candidate for the detection of He i due to its low surface gravity, large scale height, and K-type host star. However, J. Kirk et al. (2022)’s observations with Keck/NIRSPEC resulted in an upper limit of 1.25 %. This limit could be lower due to the planet’s grazing transit, and the authors noted that this tentative detection may indicate a weaker relationship between stellar high-energy activity and the formation of metastable helium. Our upper limit with HPF of 1.1 % agrees with that of J. Kirk et al. (2022).
Meanwhile, we have confirmed several previous robust and tentative detections of He i for HD189733 b, HD209458 b, WASP-52 b, WASP-69 b, and WASP-76 b. We compiled those previous spectroscopic results from the ExoAtmospheres database333https://research.iac.es/proyecto/exoatmospheres/index.php, and we compare them in Figure 6.
A He i triplet absorption ranging 0.35–0.70 % has been detected with several instruments for HD189733 b (M. Salz et al., 2018, CARMENES; G. Guilluy et al., 2020, GIANO-B; M. Zhang et al., 2022, Keck/NIRSPEC; A. Masson et al., 2024, SPIRou). However, we were only able to detect the signal (0.41 0.08 %) when using out-of-transit baseline contemporaneous to the in-transit observations. Otherwise, stellar variability over long timescales is greater than the planetary signal, which prevented us from initially detecting the He i triplet absorption. Therefore, our analysis for HD189733 b illustrates the limitations of the HET/HPF observing strategy for planetary atmospheric studies where stellar variability is a comparable signal or noise source to the signal.
HD209458 b also displays variability in its He i signal strength. L. Nortmann et al. (2018, CARMENES) and A. Masson et al. (2024, SPIRou) set their upper limits at 0.84 % and 0.5 %, respectively. However, F. J. Alonso-Floriano et al. (2019) using CARMENES found an excess absorption of 0.91 %. Using HET/HPF, we also find an excess absorption of 0.67 %, but we note that our signal’s amplitude could be attributed to correlated noise (see Sect. 3 for further details). A variable planetary signal may have contributed to this discrepancy in detections and non-detections. We need more spectroscopic transit observations to study the behavior of the He i triplet in HD209458 b’s atmosphere.
WASP-52 b has also been studied with a variety of different instruments. J. Kirk et al. (2022) and M. Fournier-Tondreau et al. (2025) reported detections of 3.44 % and 5.5 % using Keck/NIRSPEC and JWST/NIRISS, respectively. On the other hand, R. Allart et al. (2023) reported a tentative detection with an upper limit of 1.69 %. We find evidence of He i absorption, but only at a level of 1.85 0.20%, which is compatible with the upper limit reported by R. Allart et al. (2023) in the same year, but in tension with the detections. The wide spread in the reported absorptions may indicate that WASP-52 b He i signal is variable, as has been found for other exoplanets (J. P. Ninan et al., 2020; M. Zhang et al., 2022). Interestingly, the reported excess absorption levels increase with later observing dates, as seen in Figure 6. This long-term variability might be related to stellar activity W. G. Levine et al. (2024). We encourage the exoplanet community to continue observing this target in order to track its He i triplet line behavior over the next few years.
As the first exoplanet with a ground-based helium detection (L. Nortmann et al., 2018), WASP-69 b is well-studied in the current literature. Past observations with a variety of instruments detect excess absorption levels ranging from 2.35–4.31 % (3.59 0.19 %, CARMENES, L. Nortmann et al., 2018; 2.21 0.46 %, SPIRou, R. Allart et al., 2023; 2.7 0.4 %, Keck/NIRSPEC, D. Tyler et al., 2024; 3.91 0.22 %, GIANO-B, G. Guilluy et al., 2024; 3.0 0.8 %, SPIRou, A. Masson et al., 2024; 3.17 0.05 %, SPIRou, R. Allart et al., 2025). We studied WASP-69b with HET/HPF for the first time. Although we retrieved a noisy TS for our first visit, it showed a tentative absorption signal. Our second visit showed a clear excess absorption of 3.5 %, consistent with previous detections in the literature.
Transmission spectroscopic observations of WASP-76 b have reported ambiguous results. With CARMENES, N. Casasayas-Barris et al. (2021a) found a tentative absorption signal of 0.55 % but reported a conservative upper limit of 0.88 %. A. Masson et al. (2024) using SPIRou derived an upper limit of 0.4 % from the amplitude of the residuals near the He i triplet line. Our combined TS showed a clear absorption feature of 0.48 %, which is consistent with the previous observations. The discrepancy between results may be due to stellar variability and atmospheric changes that may have caused a decrease in the strength of the helium planetary signal, as A. Masson et al. (2024) suggested.
4.2 Helium circumstellar structures and stellar variability
Following Z. Zhang et al. (2023) and M. Gully-Santiago et al. (2024), we computed the He i triplet line stellar light curve to inspect the presence of circumstellar structures. Figures 13 and 14 do not show evidence of extremely extended outflows of planetary origin, like in HAT-P-32 b and HAT-P-67 b. If such outflows existed, the tails would extend for only a small fraction of the total planetary orbit and/or would not be well sampled. Alternatively, the signal could be diluted within the stellar variability.
We find that stellar helium variability is typically below or of the order of 1 %, ranging from 0.5 % to 2–4 % in some extreme cases. These results are consistent with previous work by D. M. Krolikowski et al. (2024) on He variability with HPF. They monitored the stellar He i triplet line for a sample of young host stars and found that stars older than 300 Myr seem to behave similarly to mature and old stars. These young stars exhibited intrinsic stellar variability of 1 %, ranging 0–2 %. The HET/HPF observational methodology can be used to study atmospheres of exoplanets, even those around moderately young stars.
4.3 Constraints on mass loss rates and helium absorption signals
This work presents the first analyses of the He i triplet for HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. Although their transmission spectra do not show evidence of planetary absorption, we can estimate the expected helium absorption signal based on the empirical relations from J. Orell-Miquel et al. (2024, see Sect. 6.3). However, we need to compute or estimate the X-ray (5–100 Å) and extreme ultraviolet (EUV, 100–504 Å) irradiation that the planets receive from their host star. Since there are no observations of these new targets in the XUV wavelengths444In this section, XUV refers to wavelengths of X-rays and EUV defined as between 5–504 Å., we compared their stellar properties with those of the available stars in the MUSCLES survey (K. France et al., 2016) and MEATS program (P. R. Behr et al., 2023) to find a suitable comparison star. Then, we scaled the fluxes from the proxy star to our target, accounting for their stellar radii and distances. All MUSCLES and MEATS products are available on the Barbara A. Mikulski Archive for Space Telescopes555https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html (MAST). We want to emphasize that since we scale the target’s luminosity from a proxy star, these results should be considered an order-of-magnitude approximation.
HAT-P-12 ( = 0.701 R⊙, J. D. Hartman et al., 2009) is part of the MEATS survey and its analysis is presented in P. R. Behr et al. (2023). However, the star was not observed at wavelengths below 2500 Å due to its distance ( = 143 pc, C. Soubiran et al., 2018) and low activity. P. R. Behr et al. (2023, Table 2) reports a 3 upper limit of 8.791027 erg s-1 for the X-ray luminosity . We derive 2.401028 erg s-1 using the luminosity relation between X-rays and EUV from J. Sanz-Forcada et al. (2025). Those upper limits are consistent with the luminosities from the HAT-P-12 MUSCLE model, where the X-ray and far ultraviolet fluxes are based on MUSCLES observations of HD85512. We computed 7.121026 erg s-1, and 7.811027 erg s-1. Using the from MUSCLES and the XUV and He i triplet signal relation from J. Orell-Miquel et al. (2024), we obtain an EW = 223 mÅ, consistent with our derived upper limit (21.2 mÅ). The noise level reached by our observations is below this limit. Therefore, an uncontaminated transit of HAT-P-12 b could accurately measure its mass loss. If we consider the upper limit XUV values, we estimate a much larger signal (EW 36 mÅ) which is not supported by our observations.
HAT-P-17 ( = 0.838 R⊙, A. W. Howard et al., 2012; = 92.4 pc, C. Soubiran et al., 2018) is not part of the MUSCLES survey, so we estimated its XUV flux scaling the luminosities of HD97658 ( = 0.728 R⊙, T. G. Ellis et al., 2021; = 21.1 pc, C. Koen et al., 2010) and HAT-P-26 ( = 0.79 R⊙, = 141.8 pc, J. D. Hartman et al., 2011). We computed a 2.51028 erg s-1 from the scaled HD97658 MUSCLES spectrum (R. O. P. Loyd et al., 2016), while 2.51027 erg s-1 is obtained when using the HAT-P-26 MEATS spectrum (P. R. Behr et al., 2023). According to these , we computed an expected He i triplet signal of 3–10 mÅ, which is consistent with the observational upper limit of 7.4 mÅ.
For HD118203 ( = 2.03 R⊙, G. Maciejewski et al., 2024; = 92.04 pc, C. Soubiran et al., 2018), we used as a proxy HD149026 ( = 1.34 R⊙, = 75.0 pc, B. Sato et al., 2005). We computed a 1.261028 erg s-1 from the scaled HD149026 MEATS spectrum (P. R. Behr et al., 2023). Although our observations imposed a restrictive upper limit on the presence of He i (EW 2.1 mÅ), the expected absorption is estimated to be of the same order (2 mÅ).
For TrES-1 ( = 0.807 R⊙, A. S. Bonomo et al., 2017; = 159.622 pc, C. Soubiran et al., 2018), we estimated its XUV flux using HD97658 as a proxy. We computed a 2.351028 erg s-1 from the scaled HD97658 MUSCLES spectrum. From the XUV flux and the empirical relation from J. Orell-Miquel et al. (2024), we derive an EW 91.5 mÅ. The expected EW is consistent and a bit below the upper limit imposed from the observations (11.1 mÅ).
For WASP-156 b ( = 0.826 R⊙, A. S. Polanski et al., 2024; = 121.843 pc, C. Soubiran et al., 2018), we scaled the HAT-P-26 MEATS spectrum as a proxy. We obtained a 2.461027 erg s-1. We estimated an EW absorption of 31 mÅ, well bellow the upper limit from our observations (9.5 Å). Therefore, more transit observations are required to determine more restrictive limits to WASP-156 b’s mass loss.
4.4 Helium detections and trends in the gas giant population
We combined the results on gas giants from this work, from J. Orell-Miquel et al. (2024), and from the ExoAtmospheres database666https://research.iac.es/proyecto/exoatmospheres/index.php; Accessed on 1 August 2025.. Our goal is to explore whether the He i triplet detections in the gas giant population are related to planetary and stellar properties. We obtained a list of 46 planets with 0.5 RJup and He i studies. Of these, less than one-third (14) showed evidence of He absorption.
Figure 7 presents the planet radius of gas giants with helium studies as a function of their host star’s effective temperature (), , and mass. There are no He i detections for gas giants orbiting stars with higher than 6750 K. Half of the detections are in systems with 5000 K, whose planets have 1000 K. The four planets (HAT-P-32 b, Z. Zhang et al., 2023; HAT-P-67 b, M. Gully-Santiago et al., 2024; WASP-76 b, and WASP-121 b, S. Czesla et al., 2024) showing He i outflow orbiting 6750 K stars are also among the largest sampled planets ( 1.6 RJup). These four planets have of 2000 K, but below the UHJ limit (2150 K). No helium detections have been made for exoplanets more massive than 1 MJup.
Figure 7 depicts the three different scenarios for warm, hot, and ultra-hot Jupiter-sized planets. The of the host star, used as a proxy for the spectral type, plays a major role in the detectability of helium. The detection trend for WJs may be similar to that of the general exoplanet population (J. Orell-Miquel et al., 2024). WJs mainly orbit 5000 K host stars, e.g. K-type stars. These stars have a spectral energy distribution that favors the population of the He i metastable state and its detection (A. Oklopčić, 2019). UHJs and some HJs typically orbit hotter stars than WJs, making them less suitable for populating metastable helium. Therefore, their non-detections are likely related to the stellar type of their host rather than a lack of helium atoms in their atmosphere. Inspection of evaporation through other spectral lines, such as H, would be more appropriate for UHJs orbiting A-type stars J. Orell-Miquel et al. (2024). However, the extreme irradiation of F-type stars appears to be enough to make the He i outflow detectable for low-density HJ planets (0.5 g cm-3), such as HAT-P-32 b, HAT-P-67 b, WASP-76 b, and WASP-121 b. Low-density, irradiated gas giants are expected to evaporate at a low rate until they reach a bulk density of 0.1 g cm-3. At this point, they undergo runaway mass loss, which strips away their outer atmospheric layers (D. P. Thorngren et al., 2023). This large outflow would be detectable even if the metastable population mechanism is not efficient, such as in the F-type stars environment. This evaporation process is more intense during the first few hundred Myr, but it can act at Gyr time scales as well (see animation of D. P. Thorngren et al., 2023, Fig 3). All four HJs with helium detections orbit stars older than 1 Gyr.
We also explored the stellar age parameter, but we have not found a clear relation with the He i triplet detections. There are only three gas giants with helium detection around stars younger than 1 Gyr (TOI-5398 b, M. C. D’Arpa et al., 2024; TOI-1268 b, J. Orell-Miquel et al., 2024; and WASP-52 b), but there are also less studied planets, 12 in total. Thus, the ratio of detections to non-detections in the young planet population is similar to that of the old population.
5 Conclusions
In this work, we present the results of a helium survey that aims to investigate the atmospheric mass loss in individual irradiated gas giants, but also explore population trends. We analyzed HPF high-resolution transmission spectroscopic time-series of sixteen gas giants searching for He i triplet absorption. We are able to confirm previous robust or tentative detections for HD189733 b, HD209458 b, WASP-52 b, WASP-69 b, and WASP-76 b, confirming the HET/HPF ability to precisely measure atmospheric absorption. However, HD189733 b analysis highlights the limitations of the HET/HPF observing strategy and emphasizes the importance of contemporaneous baseline observations for this method.
For the first time, we provide constraints on the helium line of HAT-P-12 b, HAT-P-17 b, HD118203 b, TrES-1 b, and WASP-156 b. For these targets, our observational upper limits are consistent with the empirical helium predictions from estimated XUV fluxes. Additionally, we do not detect significant absorption in HAT-P-3 b, WASP-11 b, WASP-80 b, WASP-127 b, and WASP-177 b, which is consistent with previous studies. HAT-P-26 b had a previous detection using narrow-band photometric observations (S. Vissapragada et al., 2022) that we are unable to reproduce with our high-resolution observations.
We compiled a list of 46 irradiated gas giants for which helium studies were conducted, combining the results presented here with those available in the literature. Using this dataset, we explore He i triplet trends within the populations of warm, hot, and ultra-hot Jupiters. We note differences in the detectability of helium among these three populations. Warm Jupiter detections seem to be more closely linked to the host star type, with all of the detections found in planets orbiting K-type stars, which are considered the most favorable stellar type according to theory (A. Oklopčić, 2019). Although there are only four hot Jupiter detections, they come from low-density planets orbiting F-type stars. To date, no helium absorption signal has been reported for ultra-hot Jupiters. Unfortunately, our results are based in limited number of explored atmospheres and further observations of irradiated gas giants are needed to increase this sample and improve our understanding of atmospheric mass loss through He i triplet observations.
Appendix A Planetary system parameters
| (a) | (b) | (b) | (c) | ||||||
| Planet | [d] | [d] | [h] | [min] | [au] | [deg] | [km s-1] | [m s-1] | [km s-1] |
| HAT-P-3 b | 2.89973815 (13) (1) | 6843.022438 (81) (1) | 2.080.01 (2) | 14.2 | 0.03878 (65) (3) | 86.310.19 (3) | 23.4 (4) (4) | 90.630.58 (2) | 145 |
| HAT-P-12 b | 3.2130598 (21) (1) | 4419.19556 (20) (1) | 2.3380.014 (1) | 18.6 | 0.0384 (3) (1) | 89.00.4 (1) | 40.0 (9) (2) | 35.81.9 (1) | 130 |
| HAT-P-17 b | 10.3385346 (9) (1) | 6569.05972 (5) (1) | 4.050.02 (2) | 32.4 | 0.0882 (14) (2) | 89.20.1 (2) | 20.13 (21) (2) | 58.80.9 (2) | 93 |
| HAT-P-26 b | 4.23450020 (64) (1) | 6901.059458 (94) (1) | 2.450.02 (2) | 11.1 | 0.0479 (6) (2) | 88.60.5 (2) | 14.57 (46) (3) | 8.51.0 (2) | 123 |
| HD118203 b | 6.1349890 (13) (1) | 8712.66178 (19) (1) | 5.6380.013 (1) | 21.50 | 0.0701 (4) (1) | 88.91.0 (1) | 29.177 (8) (1) | 222.22.1 (1) | 124 |
| HD189733 b | 2.218574944 (30) (1) | 6194.067619 (34) (1) | 1.80360.0023 (2) | 10.0 | 0.03099 (60) (1) | 85.580.06 (1) | 2.55 (16) (1) | 2056 (1) | 151 |
| HD209458 b | 2.218574944 (30) (1) | 4560.80588 (8) (2) | 2.9780.0051 (3) | 8.7 | 0.04707 (45) (1) | 86.780.07 (1) | 14.741 (2) (4) | 84.270.70 (1) | 145 (3) |
| TrES-1 b | 3.030069476 (72) (1) | 6822.891157 (63) (1) | 2.5 (2) | 18.8 | 0.03926 (58) (3) | 90.0 (3) | 20.40 (44) (4) | 115.26.2 (5) | 141 |
| WASP-11 b | 3.72247919 (18) (1) | 6646.984352 (76) (1) | 2.5560.030 (1) | 19.283 | 0.04375 (74) (2) | 89.030.34 (2) | 4.74 (54) (3) | 80.61.8 (4) | 128 |
| WASP-52 b | 1.74978117 (16) (1) | 6784.057988 (68) (1) | 1.85 (2) | 23.5 (2) | 0.0272 (3) (3) | 85.350.20 (1) | 0.48 (33) (4) | 84.33.0 (3) | 167 |
| WASP-69 b | 3.8681382 (6) (1) | 5748.83344 (18) (1) | 1.85 (1) | 23.5 (1) | 0.04525 (75) (1) | 86.710.20 (1) | 9.37215 (23) (1) | 38.12.4 (1) | 127 |
| WASP-76 b | 1.80988198 (56) (1) | 8080.6261 (4) (1) | 3.83 (1) | 23.6 (1) | 0.0330 (2) (1) | 89.620.03 (1) | 1.0733 (2) (2) | 116.01.3 (1) | 197 (3) |
| WASP-80 b | 3.06785234 (79) (1) | 6487.425006 (25) (1) | 2.1310.003 (1) | 20 (1) | 0.0344 (10) (1) | 89.020.10 (1) | 9.82 (77) (2) | 109.03.1 (1) | 122 |
| WASP-127 b | 4.17806203 (53) (1) | 6776.62124 (23) (1) | 4.3530.014 (1) | 24.9 (1) | 0.0484 (9) (1) | 87.840.33 (1) | 8.25 (89) (2) | 223 (1) | 126 |
| WASP-156 b | 3.8361623 (15) (1) | 8414.13606 (31) (1) | 2.4040.019 (1) | 9.8 (1) | 0.0458 (9) (1) | 89.10.9 (1) | 9.30 (61) (2) | 18.71.1 (1) | 130 |
| WASP-177 b | 3.071722 (1) (1) | 7994.37140 (28) (1) | 1.610.03 (1) | - | 0.03957 (58) (1) | 84.140.8 (1) | 7.1434 (41) (1) | 77.35.2 (1) | 139 |
Appendix B A note on Ca ii IRT inspection with HET/HPF
This work focuses on the He i triplet line, but the Ca ii infrared triplet (IRT; 8500 Å, 8544 Å, and 8664 Å) lines are also accessible with the HPF spectrograph. In this section, we explore HPF’s ability to detect and measure Ca ii IRT absorption signals in planetary atmospheres. In particular, we focus on WASP-76 b because N. Casasayas-Barris et al. (2021a) already reported the detection of Ca ii IRT lines combining two transits observed by the CARMENES spectrograph.
We observed two consecutive transits of WASP-76 b with HPF: visit 1 has out-of-transit baseline the night before and during the same night, and visit 2 has out-of-transit baseline one and two nights after. We find that the transmission spectra obtained using out-of-transit data from farther in time were noisier than those obtained using data from closer to the transiting night. The most reliable results are obtained from visit 1 using only the spectra taken on the same night, which is shown in Figure 8. The TS from other out-of-transit selections are strongly affected by stellar variability. Our findings are expected since the Ca ii IRT lines are used as stellar activity tracers (M. Zechmeister et al., 2018).
The 8544 Å, and 8664 Å TS exhibit 2% features, much deeper than the 0.6% and 0.35% absorptions, respectively, reported by N. Casasayas-Barris et al. (2021a). The 8500 Å TS does not have the precision to confirm the previous detection of 0.22%, and we can only to rule out absorptions greater than 1%. If the signals at 8544 Å, and 8664 Å are real, e.g., due to planetary absorption variability, then the absorption at 8500 Å should be deeper as well. However, there is no evidence for this scenario. Furthermore, the three residual maps show vertical structure at the line positions, likely originating from line core variability. The stellar Ca ii IRT lines are usually stronger and deeper than the He triplet line resulting in lower S/N at the core. The residual maps from N. Casasayas-Barris et al. (2021a, Figure 8) show residuals from the line core as well, but the planetary trace is visible at phases far from the central transit time, due to the large velocity shift of the planet. Our visit 1 with HPF only covered a fraction of the transit close to the central time, where the planetary trace intersects the core. Therefore, any planetary signal is buried in the core variability. Similar results were obtained for HAT-P-67 b. M. Gully-Santiago et al. (2024) with HPF did not detect the Ca ii IRT absorption reported by A. Bello-Arufe et al. (2023) with CARMENES.
We conclude that inspecting Ca ii IRT planetary absorption with HET/HPF is challenging due to line core variability. However, it could be feasible for planets with large velocity shifts, in-transit observations far from the central time, and out-of-transit observations on the same night.
Appendix C Reference stellar spectra and individual transmission spectra
Appendix D Stellar He triplet light curves